OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46 centimeters. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences.

terenzreignz (terenzreignz):

Well, I'm glad you're not asked to solve it :3 Anyway, how do you get the perimeter of a triangle?

OpenStudy (anonymous):

I believe that would be 2X+Y=46 centimeters and X+4=Y

OpenStudy (anonymous):

I'm not 100 percent sure on that, though.

OpenStudy (jhannybean):

You Sir are correct :)

OpenStudy (anonymous):

*lady haha but good to know that my algebra is in check

OpenStudy (anonymous):

is this good? x + x + y = 46 y = x + 4

OpenStudy (jhannybean):

sorry* ma'am:)

OpenStudy (anonymous):

thats what i had

OpenStudy (jhannybean):

Bring the common terms together charlie. x+x=?

OpenStudy (anonymous):

2x?

OpenStudy (jhannybean):

mmhmm. There you go :D

OpenStudy (anonymous):

:D :)

OpenStudy (anonymous):

I guess you were right @Sophie325 :)

OpenStudy (anonymous):

Shall I take a bow or..?

OpenStudy (anonymous):

haha

OpenStudy (jhannybean):

:P

OpenStudy (anonymous):

i got a medal? YAY!